OpenStudy (anonymous):

if x^2+y^2=16 and x^2-y^2=8 , what is the value of x^4-y^4? answers a) 1 over 2 b)4 c)8 d)128 e)256

OpenStudy (anonymous):

Use this: a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

would it be c?

OpenStudy (anonymous):

I'm afraid not. In this case a=x^2 and b=y^2. So if we know what x^2 and y^2 equals we also know what x^4-y^4 equals. So try to find x^2 and y^2 first.

OpenStudy (anonymous):

answer wud be D...that is 128

OpenStudy (anonymous):

how would it be d?

OpenStudy (anonymous):

It's correct. Have you found x^2 and y^2 yet? Tell me where you run in to problems and I'll explain.

OpenStudy (anonymous):

wouldn't it be 2 2*2 = 4 *2=8+2*2=4*2=8 so 8+8 = 16?

OpenStudy (anonymous):

Hold on, I don't see how ties in with this problem. and 2*2 is not equal to 4*2.

OpenStudy (anonymous):

so how would i do this? to get the answer

OpenStudy (anonymous):

Ok, let's find x^2. We could add or subtract the second equation to the first to get x^2. Try that.

OpenStudy (anonymous):

u find x^2 and y^2...x^2 wud be 12...y^2 wud b 4...

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

Do you understand how to find x^2 and y^2?

OpenStudy (anonymous):

yeaa , thanks(:

OpenStudy (anonymous):

you dont need to find any of the variables at all, you just need the fact that: \[x^{4}-y^{4} = (x^{2}+y^{2})(x^{2}-y^{2}) = 16*8 = 128\]